OpenStudy (accidentalaichan):

check: 2x + y = 5 x − 2y = 10

OpenStudy (accidentalaichan):

y = 5 − 2x x − 2(5 − 2x) = 10 x − 10 + 4x = 10 5x − 10 = 10 5x = 20 x = 4 2(4) + y = 5 y = −3 Is my work correct?

OpenStudy (anonymous):

right :)

OpenStudy (accidentalaichan):

Thank you :)